# On the Roles of LLMs in Planning: Embedding LLMs into Planning Graphs

### v1

@@ 4.1 Building Planning Graphs with LLMs
A planning graph $PG$ is the search space for a relaxed version of the planning problem, an intuitive framework of which is shown in Figure 5. It alternates layers of ground literals and actions. “Square” nodes at action-level $i+1$ indicate actions that might be possible to be executed in state $s_{i}$. Maintenance actions indicate dump operators that keep literals unchanged between state-levels $i$ and $i+1$. “Black circle” nodes at state-level $i$ indicate literals that might possibly be true at time $i$. Edges between state-level $i$ and action-level $i$ indicate literals in state-level $i$ are preconditions of actions in action-level $i$, while edges between action-level $i$ and state-level $i+1$ indicate literals in state-level $i+1$ are adding or deleting effects of actions in action-level $i$. The nodes in the first state-level indicate literals that are true in initial state $s_{0}$.
@@
The procedure of build the planning graph with LLMs (i.e., buildGraphLLMs) based on the given planning problem $\mathcal{P}=\langle\mathcal{R},s_{0},g,\mathcal{A}\rangle$ is as follows:
@@
    All propositions in $s_{0}$ and the negation of propositions in $\mathcal{R}-s_{0}$ are added into state-level 0.
@@
Figure 3: The prompt for consulting LLMs
@@ 4.2 Building Mutual Constraints
Due to the satisfaction of action models being relaxed, actions and/or states in action-levels or state-labels may be inconsistent, i.e., there may be some actions mutually exclusive in action-levels, or some literals mutually exclusive in state-levels. As shown in Figure 4, there are there types mutual exclusion constraints among actions.
Figure 4: Mutual exclusion of actions
Specifically, two actions at the same action-level are mutex, if they satisfy the following conditions:
@@ 4.3 Sort Action Sets with LLMs and Search Solutions
After we build the set of constraints in Step 3 of Algorithm 1, we use off-the-shelf procedure presented in (Blum & Furst, 1997) to compute candidate action sets such that there are no conflicts (i.e., satisfying the constraints $\mathcal{C}$) among actions in each action set. After that, we consult LLMs to sort the action sets by designing the prompts as shown in Figure 6, which is similar to the prompt shown in Figure 3 except the command in BLUE. After we get the sorted action sets $\mathbb{A}$, we conduct the dept-first search procedure as done in (Blum & Furst, 1997) by giving the priority of action sets based on the sorted action sets in $\mathbb{A}$.
Figure 6: The prompt for consulting LLMs to sort action sets
In Step 9 of Algorithm 1 (i.e., $expandGraphLLMs(PG)$), we expand $PG$ with the same prompt as shown in Figure 3 when selecting actions for expanding one more level of $PG$.
@@ 5.1 Experimental Setup
In this experiment, we pick four domains with different scenarios, including gripper, miconic, logistics, and movie. 10 problems are randomly selected for each domain. The specific scenarios and scale are described as follows:
@@ 5.2 Experimental Results
In the experimental framework described, we employed five distinct metrics to assess the efficacy of various methodologies: the problem-solving success rate, the cumulative count of expansion actions, the aggregate number of mutually exclusive actions, the quantity of expansion layers, and the node count for backtracking in Depth-First Search (DFS).
Problem-solving success rate. The solvability of a problem is a crucial metric in assessing planning problems. All approaches are required to generate a sequence of actions that is sufficient to solve the problem, and only if the problem can be transferred from the initial state to the goal state through this sequence of actions can the corresponding problem be considered to be successfully solved. Furthermore, we imposed a maximum time constraint of five minutes for problem resolution; any problems unsolved within this duration are deemed failures.
Expansion Layers. Expansion layers hold significant importance in the GP algorithm, as each additional expansion layer exponentially increases the complexity of DFS during backtracking. Theoretically, the GP algorithm, provided it can compute a solution, achieves the minimal number of levels and the shortest action sequence. Therefore, maintaining a minimal increase in expansion levels is pivotal and it ensures the consistent attainment of a near-optimal action sequence.
Regarding the latter four metrics, our analysis was confined to data from the GP and LLMs4Plan-GPT4 methods, primarily for two reasons. Firstly, these metrics are relevant only in scenarios where the problem is successfully solved; failed solutions do not yield countable data. Consequently, we excluded LLMs4Plan-GPT3.5 from our statistical analysis due to its comparatively lower success rate. Secondly, these metrics are inherently calculable within the GP framework alone. Hence, directly solving problems using GPT-3.5 and GPT-4 precludes the possibility of gathering this data, as these methods operate outside the GP framework.
We display the above metrics in two distinct tables. Table 1 shows the problem-solving success rates, while table 2 focuses on the last four metrics previously mentioned.
@@
gripper 0.0 0.6 0.6 0.0 0.6
miconic 0.1 0.5 0.3 0.1 0.5
logistics 0.2 0.6 0.6 0.2 0.8
movie 1.0 1.0 1.0 1.0 1.0
Table 2: LLM pruning results. In the table, each column signifies a distinct domain. It displays the statistical outcomes of the two methods across four indicators. Smaller numbers in the table denote lower computational resource usage and greater effectiveness.
Expansion Actions Mutex Actions Expansion Layers DFS Nodes
Domain GP LLMs4Plan GP LLMs4Plan GP LLMs4Plan GP LLMs4Plan
gripper 157 128 1656 1179 4 4 21321 498
miconic 181 176 792 668 5 6 2018484 11891
logistics 294 167 1726 161 5 5 1261250 59
movie 401 308 7 7 2 2 10869160 975163
5.3 Ablation Experiment
We conducted four ablation experiments to ascertain the effectiveness of forward pruning and backward sorting, detailed as follows:
@@ 5.3 Ablation Experiment
    LLMs4Plan-pruned-sorted: This method involves both forward pruning and backward sorting, corresponding to the LLMs4Plan-GPT4 approach in our experiments setup.
@@
    LLMs4Plan-pruned-unsorted: Here, we implement pruning without sorting.
@@
    LLMs4Plan-unpruned-sorted: In this approach, sorting is used, but not pruning.
@@
    LLMs4Plan-unpruned-unsorted: This method involves neither pruning nor sorting, akin to the GP method in our experimental setup.
Table 3 reveals that both pruning and sorting contribute to enhanced search efficiency, with their combination amplifying this effect. Notably, pruning appears slightly more effective than sorting. This is likely because LLM, while pruning, also organizes the remaining actions logically. In contrast, sorting may lead to minor errors due to the multitude of actions and lengthy text. The experiments also indicate that LLM tends not to prioritize empty actions in graph planning, favoring their later arrangement. This aligns with our analysis suggesting that prioritizing non-empty actions is more productive, as a layer without any action is essentially redundant.
@@
pruned-sorted pruned-unsorted unpruned-sorted unpruned-unsorted
gripper 498 1858 8317 21321
@@ 5.4 Experimental Analysis
Analyzing Table 1, GPT-3.5 only handles simpler tasks like movie, showing low success rates in other domains and thus offering limited enhancement to the GP algorithm. GPT-4 exhibits significant improvements over GPT-3.5, particularly in reasoning and the decision of long action sequences. Leveraging GPT-4’s advanced reasoning and common sense, GP demonstrates improved problem-solving success in certain domains. The observed discrepancy largely stems from GP’s timeout failures in some cases. However, with LLM-based pruning, GP successfully resolves these issues within the set time limit, thereby boosting the success rate.
Turning to the second table, we assess how the GP method, augmented by LLM, enhances its success rate. The data indicates that both the total number of expansion actions and mutually exclusive actions experience a consistent decline with LLM integration. This suggests a reduction in the computational effort required for forward expansion. Concurrently, the number of expansion layers remains relatively stable, contributing to the generation of near-optimal action sequences. During the DFS in backtracking, LLM pruning significantly cuts down the number of search nodes, thereby endowing the GP method with expedited problem-solving capabilities, including for problems that were previously unsolvable.
@@
We illustrate an example from the logistics domain in Figure 7, where we compare the number of expansion actions before and after LLM pruning. The application of LLM for pruning demonstrates significant efficacy across all layers.
@@
Figure 7: An example of action pruning. Both the GP and LLMs4Plan-GPT4 methods expanded through 10 layers. The horizontal axis represents the layer number, with lower numbers indicating proximity to the initial state and higher numbers nearing the goal state. The vertical axis shows the count of expanded actions, including both domain-specific actions and numerous empty actions, characteristic of the graph planning algorithm. This implies a high pruning ratio for genuinely effective actions. LLM prunes almost every layer of expansion actions and the data in the table also contains many empty actions.
After checking the specific generated action sequences, we found that although the solution success rates of GPT-4 and our approach were almost equivalent, the length of the action sequences generated with GPT4 alone is generally longer. In contrast, integrating GPT-4 with graph planning yields more efficient and superior action sequences.
@@
Upon analyzing examples where solutions failed, we observed that GPT4 is more prone to pruning errors at deeper expansion levels. This results in the discarding of effective actions, thereby unnecessarily increasing the expansion layers and hindering problem resolution. We attribute this to two primary factors. Firstly, as the expansion level deepens, both the predicate set and the candidate action set expand, leading to increasingly lengthy input prompts. This prolonged text can cause GPT-4 to gradually lose track of previous information, resulting in decision-making errors. Secondly, the nature of the predicate set in graph planning diverges from traditional planning’s current state representation. This discrepancy impairs LLM’s ability to accurately analyze the predicate set, leading to the erroneous elimination of effective actions. LLM lacks capacity to analyze complex predicate set combinations.
In response to GPT4’s pruning errors at deeper levels, we experimented with pruning only at shallower levels. This approach successfully avoids the earlier drawbacks, improving problem-solving success and efficiency. The key is that shallow-layer pruning with GPT-4 sidesteps the issues identified earlier. Therefore, identifying the optimal layers for LLM application in graph planning requires further research and experimentation.
@@
Our analysis of additional failure examples indicates that graph planning excels in efficiently handling numerous non-mutually exclusive actions in parallel, due to its ability to group these actions within the same layer. However, its limitation becomes apparent in scenarios requiring the execution of longer, complex actions. If a problem’s optimal solution sequences are lengthy, the planning graph must be expanded considerably deeper. Despite effective pruning, this does not resolve the issue of exponential complexity growth in backtracking DFS caused by increased depth.

### v2

@@ 4.1 Building Planning Graphs with LLMs
A planning graph $PG^{r}$ is the search space for a relaxed version of the planning problem, an intuitive framework of which is shown in Figure 2. It alternates layers of ground literals and actions. “Square” nodes at action-level $i+1$ indicate actions that might be possible to be executed in state $s_{i}$. Maintenance actions indicate dump operators that keep literals unchanged between state-levels $i$ and $i+1$. “Black circle” nodes at state-level $i$ indicate literals that might possibly be true at time $i$. Edges between state-level $i$ and action-level $i$ indicate literals in state-level $i$ are preconditions of actions in action-level $i$, while edges between action-level $i$ and state-level $i+1$ indicate literals in state-level $i+1$ are adding or deleting effects of actions in action-level $i$. The nodes in the first state-level indicate literals that are true in initial state $s_{0}$.
@@
The procedure of building the planning graph with LLMs (i.e., buildGraphLLMs) based on the given planning problem $\mathcal{P}=\langle\mathcal{R},s_{0},g,\mathcal{A}\rangle$ is as follows:
@@
    All propositions in $s_{0}$ and negation of propositions in $\mathcal{R}-s_{0}$ are added into state-level 0.
@@
Figure 3: The prompt for pruning actions
In Step 5, we use LLMs to help select actions to build the planning graph. Note that in classical graph-based planning [2], all of the actions whose preconditions are satisfied will be added into the action-level.
@@ 4.2 Building Mutual Constraints
Due to the satisfaction of action models being relaxed, actions and/or states in action-levels or state-labels may be inconsistent, i.e., there may be some actions mutually exclusive in action-levels, or some literals mutually exclusive in state-levels. As shown in Figure 4, there are three types of mutual exclusion constraints among actions. Specifically, two actions at the same action-level are mutex, if they satisfy the following conditions:
@@ 4.3 Sort Action Sets with LLMs and Search Solutions
Figure 6: The prompt for sorting action sets
After we build a set of constraints in Step 8 of Algorithm 1, we use off-the-shelf procedure presented in [2] to compute candidate action sets such that there are no conflicts (i.e., satisfying the constraints $\mathcal{C}$) among actions in each action set. After that, in Step 9, we consult LLMs to sort the action sets by designing the prompts as shown in Figure 6, which is similar to the prompt shown in Figure 3 except the command in BLUE. After we get the sorted action sets $\mathbb{A}$, in Step 10, we conduct the dept-first search procedure as done in [2] by giving the priority of action sets based on the sorted action sets in $\mathbb{A}$.
@@ 5.1 Experimental Setup
In the experiment, we evaluate LLMs4Plan in ten planning domains with different scenarios, including gripper, miconic, logistics, movie, blocks, satellite, zenotravel, driverlog, woodworking and openstacks. Ten problems are randomly selected for each domain. The specific scenarios and sizes are described in Appendix A.1.
@@
5.2 Experimental Metrics
In the experimental framework described, we employed three distinct metrics to assess the efficacy of various methodologies: the problem-solving success rate, the cumulative count of expansion actions and the node count for backtracking in Depth-First Search (DFS).
Problem-solving success rate. The solvability of a problem is a crucial metric in assessing planning problems. All approaches are required to generate a sequence of actions that is sufficient to solve the problem, and only if the problem can be transferred from the initial state to the goal state through this sequence of actions can the corresponding problem be considered to be successfully solved. Furthermore, we have established an upper bound on the depth of the problem-solving process. The optimal length of the action sequence for the test problem is known to us. Should the solution obtained surpass this predetermined depth, it signifies the inability of the method to successfully ascertain the optimal action path for this particular problem. Setting an upper bound on the depth of the problem-solving process serves the purpose of not only requiring the planner to solve problems but also demanding that it does so more efficiently. This ensures that the output action sequences are more concise and accurate, minimizing the occurrence of redundant actions.
Regarding the number of nodes for backtracking DFS, our analysis was confined to data from the GP and LLMs4Plan-GPT4 methods, primarily for two reasons. Firstly, the metric is relevant only in scenarios where the problem is successfully solved; failed solutions do not yield countable data. Consequently, we excluded LLMs4Plan-GPT3.5 from our statistical analysis due to its comparatively lower success rate. Secondly, these metrics are inherently calculable within the GP framework alone. Hence, directly solving problems using GPT-3.5 and GPT-4 precludes the possibility of gathering this data, as these methods operate outside the GP framework.
@@ 5.3 Experimental Results
We present the success rates in Table 1, depict the pruning effects of action expansion for LLM on GP in Figure 7, and showcase experimental results in Table 2 comparing our approach to traditional GP algorithms in terms of the number of nodes for backtracking DFS metrics, along with relevant ablation studies.
@@
gripper 0.00 0.60 0.70 0.00 1.00
miconic 0.10 0.50 0.60 0.10 1.00
logistics 0.20 0.60 0.60 0.20 1.00
movie 1.00 1.00 1.00 1.00 1.00
blocks 0.10 0.70 0.60 0.30 1.00
satellite 0.00 0.50 0.90 0.10 1.00
zenotravel 0.20 0.60 0.90 0.20 1.00
driverlog 0.00 0.10 0.90 0.20 1.00
woodworking 0.90 0.90 0.70 1.00 1.00
openstacks 0.10 0.20 1.00 0.20 1.00
Ablation Experiment: We conducted four ablation experiments to ascertain the effectiveness of forward pruning and backward sorting, detailed as follows:
@@
    LLMs4Plan: This method involves both forward pruning and backward sorting.
@@
    LLMs4Plan-unsorted: Here, we implement pruning without sorting.
@@
    LLMs4Plan-unpruned: In this approach, sorting is used, but not pruning.
@@
    GP: This method involves neither pruning nor sorting.
@@
LLMs4Plan LLMs4Plan-unsorted LLMs4Plan-unpruned GP
gripper 6839 11294 4850376 8486698
@@
blocks 4572 7272 83129 1205223
satellite 46619 94811 67167049 88779785
zenotravel 1548 12166 839527 2259283
driverlog 574 1916 58311 1579486
woodworking 49 3924 48553 114502
openstacks 107 409 13577 24267
@@ 5.4 Experimental Analysis
Analysis of the success rate of planning: From Table 1, several conclusions can be drawn. GPT3.5 exhibits competence primarily in resolving simple problems with short action sequence lengths, such as in the movie domain, while its success rates are notably low in other domains. Consequently, it struggles to enhance the capabilities of GP algorithms. Conversely, GPT4 demonstrates substantial improvements in abilities compared to GPT3.5, particularly in reasoning skills and decision-making involving long action sequences. With the enhanced reasoning and commonsense capabilities of GPT4, GP shows an enhanced success rate in certain domains. We observe instances of failure in GP, attributed to the inclusion of partially corrupt data during testing. Specifically, we introduce a proportion of corrupted data by randomly removing action preconditions and effects propositions from domain files. These instances have varying impacts across different domains, particularly affecting traditional GP algorithms reliant on the completeness of domain files. We detail the influence of random removal on success rates in Table 3. For each layer of GP expansion, the provision of action preconditions and effects propositions is essential. However, in the case of LLM-augmented GP algorithms, LLMs4Plan is capable of making rational action decisions even in the presence of missing action propositions, thereby aiding in the completion of current planning tasks. Consequently, our algorithm exhibits greater robustness in terms of success rates compared to traditional GP approaches. The specific settings for the robustness experiments are detailed more extensively in section A.2 of the appendix.
Table 3: This table illustrates experiments on the robustness of missing action predicates. In the table, each row corresponds to a distinct domain. Each column in the table represents the proportion of predicates we removed. A higher proportion indicates a greater amount of missing information, posing increased difficulty for the planner to solve the problem. The values in the table represent the success rates of GP in solving the problems. In the majority of domains, as the proportion of deleted predicates increases, the success rate of GP planning decreases. Overall, this indicates that GP exhibits poor robustness to missing action predicates.
10% 20% 30% 40% 50%
gripper 0.40 0.87 0.73 0.80 0.67
miconic 0.60 0.60 0.60 0.80 0.40
logistics 0.20 0.80 0.80 0.67 0.60
movie 1.00 1.00 1.00 1.00 0.60
blocks 1.00 0.26 0.07 0.27 0.47
satellite 0.60 0.73 1.00 0.86 0.87
zenotravel 0.93 0.93 1.00 0.93 0.80
driverlog 0.80 1.00 1.00 1.00 1.00
woodworking 1.00 0.73 0.40 0.27 0.40
openstacks 1.00 1.00 1.00 1.00 1.00
Upon examining the generated action sequences, we observed that although GPT4 achieves a certain level of success in solving problems, the action sequences it produces tend to be longer compared to those generated by GP alone. By integrating GPT4 with graph planning, LLMs4Plan can effectively generate more optimal action sequences.
Analysis of search efficiency: Besides planning success rates, our method significantly improves search efficiency compared to GP algorithms. This enhancement is evident from Table 2, where, among problems with successful planning outputs, we drastically reduce the cost of search nodes, achieving an exponential level of optimization. So, how does the LLM-augmented GP method enhance search efficiency?
Through in-depth analysis of experimental cases, we identify two main aspects of optimization:
  1. 1.
    During forward expansion, LLM efficiently and effectively prunes the expansion actions, leading to varying degrees of stable reduction in the total number of expanded actions and mutually exclusive actions. Consequently, the computational load of forward expansion decreases correspondingly.
  2. 2.
    During the depth-first search backtracking process, LLM prioritizes searching closer to the set of planning solutions, accelerating the attainment of planning solutions and saving time by avoiding ineffective searches.
@@
We provide an example from the ’logistics’ domain to illustrate our analysis in Figure 7, where we compare the number of expansion actions before and after LLM pruning. The application of LLM for pruning demonstrates significant efficacy across all layers. In addition, we have provided further experimental results and analysis on the total number of mutually exclusive actions and expanded actions in section A.3 of the supplementary materials.
@@
Figure 7: An example of action pruning. Both the GP and LLMs4Plan-GPT4 methods expanded through 10 layers. The horizontal axis represents the layer number, with lower numbers indicating proximity to the initial state and higher numbers nearing the goal state. The vertical axis shows the count of expanded actions, including both domain-specific actions and numerous empty actions, characteristic of the graph planning algorithm. This implies a high pruning ratio for genuinely effective actions. LLMs prune almost every layer of expansion actions and the data in the table also contains many empty actions.
For pruning, the greatest risk is removing necessary actions, rendering the problem unsolvable. In our experiments, we observed instances where LLM prunes crucial actions in certain layers, resulting in the inability to obtain effective solutions. However, we introduced pruning probabilities to ensure algorithm completeness. Experimental results demonstrate that although the process of correcting LLM’s erroneous pruning behavior through pruning probabilities may introduce additional expansion and search steps, the cumulative cost of these search steps remains significantly lower than the cost of solely using GP algorithms to solve problems. The results presented in Table 2 compare the outcomes of our method LLMs4Plan, which ensures completeness, with those of GP algorithms.
Analysis of ablation experiments: Table 2 reveals that both pruning and sorting contribute to enhanced search efficiency, with their combination amplifying this effect. Notably, pruning appears slightly more effective than sorting. This is likely because LLM, while pruning, also organizes the remaining actions logically. In contrast, sorting may lead to minor errors due to the multitude of actions and lengthy text. In this regard, we require further optimization of natural language processing techniques tailored for handling extremely long texts to enhance our framework’s capability in solving more complex problems. The experiments also indicate that LLM tends not to prioritize empty actions in graph planning, favoring their later arrangement. This aligns with our analysis suggesting that prioritizing non-empty actions is more productive, as a layer without any action is essentially redundant.
@@
Analysis of the advantages and disadvantages of LLMs4Plan: Upon analyzing examples where solutions failed, we observed that GPT4 is more prone to pruning errors at deeper expansion levels. This results in the discarding of effective actions, thereby unnecessarily increasing the expansion layers and hindering problem resolution. We attribute this to two primary factors. Firstly, as the expansion level deepens, both the predicate set and the candidate action set expand, leading to increasingly lengthy input prompts. This prolonged text can cause GPT-4 to gradually lose track of previous information, resulting in decision-making errors. Secondly, the nature of the predicate set in graph planning diverges from traditional planning’s current state representation. This discrepancy impairs LLM’s ability to accurately analyze the predicate set, leading to the erroneous elimination of effective actions. LLM lacks capacity to analyze complex predicate set combinations.
@@
Our analysis of additional failure examples indicates that graph planning excels in efficiently handling numerous non-mutually exclusive actions in parallel, due to its ability to group these actions within the same layer. However, its limitation becomes apparent in scenarios requiring the execution of highly complex and extremely long action sequences. If a problem’s optimal solution sequences are lengthy, the planning graph must be expanded considerably deeper. Despite effective pruning, this does not resolve the issue of exponential complexity growth in backtracking DFS caused by increased depth.